OpenStudy (anonymous):

F(x)=-4x-1,H(x)=-x+5 Find (HoF)(5)

OpenStudy (amistre64):

substitute H into F like this: -4(H(x)) -1 -4(-x+5) -1 ; x=5

OpenStudy (amistre64):

-4(-(5)+5) - 1 -4(0) - 1 = -1 (HoF)(5) = -1